OpenStudy (anonymous):

Solve 2x + 3y = 18 for y

jimthompson5910 (jim_thompson5910):

\[\Large 2x + 3y = 18 \] \[\Large 3y = -2x+18 \] \[\Large y = -\frac{2}{3}x+\frac{18}{3} \] \[\Large y = -\frac{2}{3}x+6 \]